A mixture of particular gas and Nitrogen gives an average molecular weight of 30 and
similar mixture (same mole fractions) of same gas with carbon dioxide gives an average
molecular weight of 38. Identify the gas based on the calculations of the molecular weight of
unknown gas?
1
Expert's answer
2020-07-01T07:08:33-0400

Solution.

"n-" number of moles;

"X -" molecular mass unknown gas;

"\dfrac{nX+nN_2}{n+n}=30;"


"\dfrac{nX+nCO_2}{n+n}=38;"

"\dfrac{nX+28n}{2n}=30;"

"\dfrac{nX+44n}{2n}=38;"

"\dfrac{nX+28n}{30}=\dfrac{nX+44n}{38};"

"8nX=256n;"

"X=32;"

Unknown gas - "O_2" ;

Answer: Unknown gas - O2.